OpenStudy (anonymous):

sin 30 = 1/2 and cos 30 - sqrt(30)/2 true or false

OpenStudy (campbell_st):

so can I just check is the question \[\cos(30) = \frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

nevermind i figured it out, thanks anyway